# GWQ: Gradient-Aware Weight Quantization for Large Language Models

### v2

@@ Abstract
Large language models (LLMs) show impressive performance in solving complex language tasks. However, its large number of parameters present significant challenges for the deployment and application of the model on edge devices. Compressing large language models to low bits can enable them to run on resource-constrained devices, often leading to performance degradation. To address this problem, we propose gradient-aware weight quantization (GWQ), the first quantization approach for low-bit weight quantization that leverages gradients to localize outliers, requiring only a minimal amount of calibration data for outlier detection. GWQ retains the weights corresponding to the top 1% outliers preferentially at FP16 precision, while the remaining non-outlier weights are stored in a low-bit format. GWQ found experimentally that utilizing the sensitive weights in the gradient localization model is more scientific compared to utilizing the sensitive weights in the Hessian matrix localization model. Compared to current quantization methods, GWQ can be applied to multiple language models and achieves lower PPL on the WikiText2 and C4 dataset. In the zero-shot task, GWQ quantized models have higher accuracy compared to other quantization methods. GWQ is also suitable for multimodal model quantization, and the quantized Qwen-VL family model is more accurate than other methods. Zero-shot target detection task dataset RefCOCO outperforms the current stat-of-the-arts method SPQR. GWQ achieves 1.2$\times$ inference speedup in comparison to the original model, and effectively reduces the inference memory.
GWQ: Gradient-Aware Weight Quantization for Large Language Models
Yihua Shao1,2, Siyu Liang3, Zijian Ling3, Minxi Yan3, Haiyang Liu4, Siyu Chen2, Ziyang Yan5, Chenyu Zhang5, Haotong Qin6*, Michele Magno6, Yang Yang2, Zhen Lei2, Yan Wang3, Jingcai Guo7, Ling Shao8, Hao Tang1* 1PKU  2CASIA  3THU  4USTB  5UNITN  6ETHz  7PolyU  8UCAS thanks: The first two authors contributed equally. correspondence authors: qinhaotong@gmail.com, haotang@pku.edu.cn
@@ 1 Introduction
Large language models (LLMs) (Touvron et al., 2023a; Achiam et al., 2023; Almazrouei et al., 2023; Touvron et al., 2023b) based on Transformer (Vaswani, 2017) have demonstrated their outstanding capabilities in scenarios such as complex question and answer, language modeling (Brown, 2020), and so on. It can often generate more accurate answers with different inputs. Its ability to handle complex linguistic tasks is due to its huge pre-trained corpus (Kaplan et al., 2020) and an astronomical number of parameters (Chowdhery et al., 2023). Its outstanding performance is rapidly generating huge applications in people’s lives. However, the huge memory consumption generated by its pre-training and inference phases leads to great difficulties in its application and deployment (Li et al., 2024a). Therefore, when deploying large language models, there are often huge GPU clusters to support the inference of the models (Xu et al., 2024).
In order to make large language models adaptable to resource-constrained edge devices (Li et al., 2024b; Huang et al., 2024), model compression (Ma et al., 2023; Gu et al., 2023) has become a common means to reduce the computational memory of models. Among them, post-training large language model quantization (Dettmers and Zettlemoyer, 2023) is a commonly used method to reduce the model inference overhead. Most current quantization methods compress the model to 3 or 4 bits (Dettmers et al., 2023; Frantar et al., 2022a), while trying to ensure that the model has as little loss of accuracy and performance as possible. Post-training quantization of pre-trained models reduces model inference overhead by allowing the number of parameter bits to decrease. This approach produces a loss of accuracy that affects the performance of the model in use. Therefore, to reduce the size of the model while ensuring its performance, it is necessary to design effective quantization algorithms (Lin et al., 2024). With effective quantization algorithms, researchers can make the performance of low-bit models infinitely close to that of 16-bit models.
OBQ (Frantar and Alistarh, 2022) indicates that when the pre-trained model has fully converged, the model should exhibit zero gradients. However, our experiments reveal that the LLMs still generate gradients in response to different text inputs, indicating that the model has not completely converged and remains sensitive to the input. Inspired by this phenomenon, we propose a gradient-aware post-training weight-only quantization method called GWQ. GWQ is the first post-training quantization approach to utilize gradients to locate outliers in pre-trained models. It introduces two key innovations. First, GWQ selects 1% of the weights as outliers by looking at the magnitude of the gradient for the response to the calibration set, preserving these outliers in FP16 precision while quantizing the remaining 99% of the weights to 4 bits or 3 bits.
@@
    GWQ discovered that employing the first-order gradient to search for and locate sensitive weights is more rational than using the Hessian matrix for localization, and the model loss after quantization is also lower. Therefore, using a first-order gradient to locate outliers is more reasonable.
@@
    GWQ is the first accurate first-order gradient-aware post-training weight quantization method for pre-trained LLMs, requiring only a minimal quantity of calibration data to identify outliers efficiently.
@@
    GWQ outshines the current state-of-the-arts method SPQR on the wikitext and C4 datasets. The majority of zero-shot detection tasks in the multimodal dataset RefCOCO have reached the state-of-the-arts. Meanwhile, the quantified model has achieved a 1.2$\times$ acceleration compared to the original model and utilized less memory during inference.
@@ 2 Related Work
Post-training Quantization. Post-training quantization (PTQ) applies mainly to visual models (Gholami et al., 2022). For example, AdaRound (Nagel et al., 2020), AdaQuant (Hubara et al., 2020), BitSplit (Wang et al., 2024b) can be used for models with a small number of parameters. When models have a large number of parameters, they cause a significant loss of accuracy, and the reduction in memory overhead of the quantized model compared to the original model is not particularly significant. Most current quantization methods for models are improved by Optimal Brain Quantization (OBQ) (Frantar and Alistarh, 2022). OBQ is improved by the Optimal Brain Surgeon (OBS) (Hassibi et al., 1993; Singh and Alistarh, 2020; Frantar et al., 2021), both OBS and OBQ default to a model whose response to input is 0 at the end of training, but the OBQ method applies the idea of weight pruning to model quantization by quantifying the weights that have the least impact on the network and compensating for them with the remaining weights. Some current PTQ methods perform outlier search by reinforcement learning (Wang et al., 2024b) to retrieve the best outlier by an objective function on a calibration set with a very large sample size. However, the training process of reinforcement learning is not robust and is difficult to converge compared to other PTQ methods, making the calibration process very inefficient.
@@
Large Language Model Quantization. Currently, large language model quantization is mainly divided into quantization-based training (Liu et al., 2023; Bondarenko et al., 2024; Zhu et al., 2023) and post-training quantization. Quantization during training is very difficult to apply due to the large amount of computational resources and overhead required, as well as the large amount of training data. The remaining methods such as ZeroQuant (Yao et al., 2022) utilize knowledge distillation to train one transformer layer at a time instead of training all transformer layers directly. However, poor quality of certain data can also affect the performance of the model after quantization (Wang et al., 2024a). Post-training quantization requires less data compared to quantization-aware training, and only a few thousand or even a few hundred calibration sets are needed to complete the retrieval of sensitive weights for the model. This greatly reduces the data cost and the computational cost of model quantization. Both GPTQ (Frantar et al., 2022a) and SPQR (Dettmers et al., 2023), quantization methods derived from the OBQ method, have shown superior performance in large language models. There are also methods such as AWQ (Lin et al., 2024), LLM.int8() (Dettmers et al., 2022) that determine the outlier of the model by searching for the activation of the model, and then quantify it with the corresponding method. However, the method of searching for outliers by activation has poor interpretability.
@@
Table 1: Gradient quantization of non-outliers with Hessian in comparison to the C4 dataset (Raffel et al., 2020). We localize outliers by gradient and Hessian on the llama-2-hf series with the llama-3-hf model and quantify non-outliers with RTN (Nagel et al., 2020) to 4 bits.
@@
Llama-2-7b-hf 6.97 13.4328 9.3408 7.1161 13.4358 9.3442 7.1187
Llama-2-13b-hf 6.47 11.5711 7.7271 6.5120 11.5755 7.7299 6.5142
Llama-3-8b-hf 8.99 18.4334 14.4928 11.5442 18.4390 14.4942 11.5483
@@ 3 The Proposed Method
The outlier search method based on Hessian matrix (Frantar et al., 2022a; Dettmers et al., 2023) is built upon the OBQ approach (Frantar et al., 2022b). The OBQ method retrieves the parameters that have the least impact on the model parameters by using the following equation:
@@
where $g_{i}=\frac{\partial E}{\partial w_{i}}$ is the first order gradient. In the OBQ framework, it is suggested that once a pre-trained model has fully converged, its gradients $g_{i}=\frac{\partial E}{\partial w_{i}}$ should ideally approach zero. However, our experiments demonstrate that LLMs continue to generate gradients in response to various text inputs. Bondarenko et al. (2023) denotes that the occurrence of outliers in large models arises from attention heads attempting to avoid updating hidden states. During this process, the softmax function magnifies the formation of strong outliers. Building on this observation, we hypothesize that when a well-trained LLM computes gradients for text input, it often focuses these gradients on irrelevant outliers as a mechanism to prevent hidden state updates. As shown in Fig. 1, compared to SPQR, GWQ demonstrates a sparser allocation of outliers. Specifically, the outliers identified are distributed either per-channel or per-row, which contrasts with the more dense and uniform distribution of outliers observed in SPQR. This sparsity in GWQ results in a more reasonable allocation, as it does not require to concentrate on the compensation of sensitive weight errors across every layer output. As shown in Tab. 1, with different proportions of outliers, the model using Hessian matrix search and quantized using RTN (Nagel et al., 2020) performs worse than the model using gradient search and quantized using RTN on the C4 dataset (Raffel et al., 2020), so it can be learned that it is more reasonable to use the gradient to do outliers localization.
Figure 1: Comparison of GWQ and SPQR outliers. Taking query (Q) in Self-Attention as an example, the outliers localized by GWQ (left figure) are more sparse compared to SPQR (right figure), and the row distribution is more pronounced, which is more friendly to subsequent quantization.
@@ 3.1 Sensitive Weight Location
GWQ captures the absolute gradients with respect to LLM, acquired through back-propagation, the detail is shown in Alg. 1. We disabled bias in all selected linear layers, since the bias term does not get multiplied by inputs in the same way as weights do. This can cause the weight gradients to be unrepresentative of the actual importance of the weights, leading to a disproportionate influence on the MSE loss calculation. As it shown in Fig. 2, the calibration dataset is represented as $D_{c}$, the weights of LLM as $W$, the loss function of LLM as $\mathcal{L}=(W;D_{c})$. The gradients $g$ to the LLM is:
$g=\bigtriangledown_{W}\mathcal{L}=\bigtriangledown_{W}\mathcal{L}(W;D_{c}).$ (2)
@@
In order to search for quantitatively optimal outliers, we define the quantization process as an optimization problem whose process can be expressed as follows:
$Q\left(W\right)=arg\min_{Q}\left\|W-W_{Q}\right\|_{2}^{2}.$ (4)
Therefore the final weight optimization formula can be expressed as:
@@
Algorithm 1 GWQ Gradient Accumulation
func get_gradient($M,X$)
1:a:=$a:=\emptyset$triangleright Set of accumulated gradients
[…]
3:  $o$ := M($X_{i}[0]$) $\triangleright$ model output
4:  $l$ := CrossEntropy($o$,$X_{i}[1]$)
5:  $l$.backward()
6:  $a$ := accum_grad($M$)
7:  $M$.zero_grad()
[…]
9:  $a_{j}:=\frac{a_{j}}{\text{dim}(X)}$
10:  $M_{g}:=a_{j}$triangleright update the model gradients
11:return $M$
func accum_grad($M,a$)
1:$q$ := [self_attn, mlp] //modules to be quantized
2:for $i=1,\dots,\text{dim}(a)$ do
3:  $a_{i}:=a_{i}+M_{a_{i}}$triangleright accumulate the gradients of each module in the model
4:return $a$
@@ 3.2 Gradient Aware Quantization
To quantize the LLM weight $W$, we first compute the scale $s$ and the zero point $z$, which are calculated as depicted in Eq. (6) and (7), where $\beta$ represents the group size (set to 16 in this case) and $b$ is the bit-width for quantization. The detail is shown in Alg. 2.
@@
We select the weight by absolute values of the weight gradients $|W_{grad}|$, the largest 1% of them are filtered as outliers, and the rest are used as quantization weights. This process can be expressed as:
@@
The quantization and dequantization operations are carried out per channel within each group to assess the quantization error. The formulas for quantization and dequantization are provided below (Eq. (9) and (10)), and any identified outliers are excluded from the error computation (Eq. (8)):
$q=round(\frac{W}{s})+z.$ (9)
The quantized weight $Q(W)$ can be denoted as:
@@
Algorithm 2 GWQ Quantization Algorithm
func GWQ_Quantize($M,X,b,\beta,p$)
Input: M – the model to be quantized
$X\in\mathbb{R}^{n\times d}$ – calibration data
$b$ – base number of quantization bits
$\beta$ – quantization group sizes
$p$ – global outlier percentage
1:$M$ :=get_gradient($M,X$)
2:$L$ :=$M$.layers
3:$n$ := dim(layers)
4:$l_{i_{s}}$ := $L_{i}$.sublayers
5:for $i=1,\dots,\text{dim}(L)$ do
6:for $j=1,\dots,\text{dim}(l_{i_{s}})$ do
7:  $W:=$
@@
10:  $l_{i_{s}}.W:=W$
11:return M
func GWQ_Sublayer_Quantize($W,W_{g},b,\beta,p_{s}$)
Input: $W\in\mathbb{R}^{m\times n}$ – weight matrix
$W_{g}\in\mathbb{R}^{m\times n}$ – gradient of weight matrix
$p_{s}$ – sublayer outlier percentage
1:$Q$ := int_matrix(m,n) //temporary quantized weight
2:$O$ := $\mathbb{I}\left(|W_{g}|\geq\text{quantile}(\alpha_{1})\right)$
3:for $i=1,\beta 1,2\beta 1,\dots,n$ do
4:  $\vec{s},\vec{z}$ := fit_quantizer($W_{:,i:i+\beta}$, $O_{:,i:i+\beta}$)
5:  for $j=i,\dots,i+\beta$ do
6:   $Q_{:,j}$ := quantize($W_{:,j},\vec{s},\vec{z}$)
7:   $\vec{w_{q}}$ := dequantize($Q_{:,j},\vec{s},\vec{z}$)
8:   $W_{:,j}:=\vec{w_{q}}\dot{(}1-O_{:,j}))$
9:return $W$
func fit_quantizer($M,\beta$)
1:$m_{i}:=\text{flatten}(M)$
2:$s_{i}:=\text{vectors}(O)$
3:for $i=1,\beta_{1},\beta_{2},\dots,\text{dim}(m)$ do
4:  $s_{i}:=\frac{\text{max}(m)-\text{min}(m)}{\text{dim}(m)}$
5:  $z_{i}:=\min(\frac{m-s_{i}}{s_{i}})$
6:return $s,z$
func quantize($W,s,z$)
1:return $\lfloor\frac{W}{s}+z+0.5\rfloor$
func dequantize($Q,s,z$)
1:return $s\cdot(Q-z)$
4 Experiments
Table 2: Perplexity of Llama-2-hf family and Llama-3-8B-hf on language task datasets. GWQ outperforms other quantization methods, especially when the average bit-width of the model is significantly lower.
Llama-2-7B-hf Llama2-13B-hf Llama-3-8B-hf

### v3

# Gradient-Aware Weight Quantization for Large Language Models

@@ Abstract
Large language models (LLMs) show impressive performance in solving complex language tasks. However, its large number of parameters presents significant challenges for the deployment. So, compressing LLMs to low bits can enable to deploy on resource-constrained devices. To address this problem, we propose gradient-aware weight quantization (GWQ), the first quantization approach for low-bit weight quantization that leverages gradients to localize outliers, requiring only a minimal amount of calibration data for outlier detection. GWQ retains the top 1% outliers preferentially at FP16 precision, while the remaining non-outlier weights are stored in a low-bit. We widely evaluate GWQ on different task include language modeling, grounding detection, massive multitask language understanding and vision-language question and answering. Results show that models quantified by GWQ performs better than other quantization method. During quantization process, GWQ only need one calibration set to realize effective quant. Also, GWQ achieves 1.2$\times$ inference speedup in comparison to the original model and effectively reduces the inference memory.
Gradient-Aware Weight Quantization for Large Language Models
Yihua Shao1,2, Yan Gu3, Siyu Chen3, Haiyang Liu3, Zijian Ling4, Minxi Yan2, Ziyang Yan5, Chenyu Zhang5, Michele Magno6, Haotong Qin6*, Yan Wang3, Jingcai Guo7, Ling Shao1* Hao Tang1* 1PKU  2CASIA  3THU  4USTB  5UNITN  6ETHz  7PolyU  8UCAS thanks: The first two authors contributed equally. correspondence authors: qinhaotong@gmail.com, haotang@pku.edu.cn
@@ 1 Introduction
Large language models (LLMs) (Touvron et al., 2023a; Achiam et al., 2023; Almazrouei et al., 2023; Touvron et al., 2023b) based on Transformer (Vaswani, 2017) have demonstrated their outstanding capabilities in scenarios such as language modeling (Brown, 2020). Its ability to handle complex linguistic tasks is due to its large pre-trained data (Kaplan et al., 2020; Yan et al., 2024, 2025) and an astronomical number of parameters (Chowdhery et al., 2023). However, the huge memory consumption leads to great difficulties in its application and deployment (Li et al., 2024a; Remondino et al., 2023). Therefore, when deploying large language models, there are often huge GPU clusters to support the inference of the models (Xu et al., 2024).
In order to apply LLMs to resource-constrained edge devices (Li et al., 2024b; Huang et al., 2024), model compression (Ma et al., 2023; Gu et al., 2023) has become common means to reduce the computational memory. Among them, post-training quantinazation of LLMs (Dettmers and Zettlemoyer, 2023) is a widely used method. Most current quantization methods target to compress modela to 3 or 4 bits (Dettmers et al., 2023; Frantar et al., 2022a), while ensuring models has less loss of performance. Therefore, reducing the size of the model while ensuring its performance is necessary by design effective quantization algorithms (Lin et al., 2024b). With effective quantization algorithms, researchers can make low-bit models’ performance close to that of 16-bit models.
OBQ (Frantar and Alistarh, 2022) indicates that when the pre-trained model has converged, the model should exhibit zero gradients. However, our experiments reveal that the LLMs still generate gradients in response to different text inputs. Inspired by this phenomenon, we propose a gradient-aware post-training weight-only quantization method called GWQ. GWQ is the first post-training quantization approach to utilize gradients to locate outliers in pre-trained models. As same as other quantization method Lin et al. (2024a); Dettmers et al. (2023), GWQ selects 1% sensitive weights as outliers by searching the gradient for the response to the calibration set, preserving these outliers in FP16 precision while quanting the remaining 99% of the weights to 4 bits or 3 bits.
@@
    GWQ discovered that locating sensitive weights by first-order gradient is more rational than Hessian matrix.
@@
    GWQ is the first accurate first-order gradient-aware post-training weight quantization method for pre-trained LLMs, requiring only one of calibration data to locating outliers.
@@
    GWQ achieves state-of-the-art (SOTA) in language modeling, grounding detection, massive multitask language understanding and vision-language question and answering tasks in different LLMs. Meanwhile, the quantified models have achieved 1.2$\times$ acceleration compared to the original model and cost less memory during inference.
@@ 2 Related Work
Post-training Quantization. Post-training quantization (PTQ) applies mainly to visual models (Gholami et al., 2022). For example, AdaRound (Nagel et al., 2020), AdaQuant (Hubara et al., 2020), and BitSplit (Wang et al., 2024b) can be used for models with less parameters. When models have a large number of parameters, they cause a significant loss but the reduction in memory is not significant. Most quantization methods are developed by Optimal Brain Quantization (OBQ) (Frantar and Alistarh, 2022). OBQ is developed by the Optimal Brain Surgeon (OBS) (Hassibi et al., 1993; Singh and Alistarh, 2020; Frantar et al., 2021), which default a model response to input is 0 at the end of training. But the OBQ applies weight pruning to model quantization by quantifying the weights that have the less impact on the network. Some PTQ methods locating outlier by reinforcement learning (Wang et al., 2024b) to retrieve the best outlier on a calibration set with a very large sample size. However, the reinforcement learning is not robust and is difficult to converge compared to other PTQ methods, making the calibration process inefficient.
@@
Large Language Model Quantization. Currently, large language model quantization is mainly divided into quantization-based training (Liu et al., 2023; Bondarenko et al., 2024; Zhu et al., 2023; Shao et al., 2025b, a) and post-training quantization. Quantization during training is very difficult to apply due to the large amount of computational resources and overhead required, as well as the large amount of training data. The remaining methods such as ZeroQuant (Yao et al., 2022) utilize knowledge distillation to train one transformer layer at a time instead of training all transformer layers directly. However, poor quality of certain data can also affect the performance of the model after quantization (Wang et al., 2024a). Post-training quantization requires less data compared to quantization-aware training, and only a few thousand or even a few hundred calibration sets are needed to complete the retrieval of sensitive weights for the model. This greatly reduces the data cost and the computational cost of model quantization. Both GPTQ (Frantar et al., 2022a) and SPQR (Dettmers et al., 2023), quantization methods derived from the OBQ method, have shown superior performance in large language models. There are also methods such as AWQ (Lin et al., 2024b), LLM.int8() (Dettmers et al., 2022) that determine the outlier of the model by searching for the activation of the model, and then quantify it with the corresponding method. However, the method of searching for outliers by activation has poor interpretability.
Figure 1: Compared of GWQ and SPQR. Compared to SPQR, the outliers searched by GWQ are much sparser and easier to quantization.
@@
Table 1: Gradient quantization of non-outliers with Hessian in comparison on C4 dataset (Raffel et al., 2020). We localize outliers by gradient and Hessian and quant the non-outliers with RTN (Nagel et al., 2020) to 4 bits.
@@
Llama-2-7b-hf 6.9759 20.5896 16.8329 15.4319 24.0418 20.2053 18.5902
Llama-2-13b-hf 6.4734 20.3549 16.6689 15.1590 23.9322 20.1040 18.3405
Llama-3-8b-hf 8.9990 22.4648 18.9843 17.5901 22.4423 22.9650 19.3283
@@ 3 The Proposed Method
The outlier search method based on Hessian matrix (Frantar et al., 2022a; Dettmers et al., 2023) is built upon the OBQ approach (Frantar et al., 2022b). The OBQ method retrieves the parameters that have the least impact on the model parameters could be expressed as Eq. (1):
@@
where $g_{i}=\frac{\partial E}{\partial w_{i}}$ is the first order gradient. In the OBQ framework, it is suggested that once a pre-trained model has fully converged, its gradients $g_{i}=\frac{\partial E}{\partial w_{i}}$ should ideally approach zero.
However, our experiments demonstrate that LLMs continue to generate gradients in response to various text inputs. Bondarenko et al. (2023) denotes that the occurrence of outliers in large models arises from attention heads attempting to avoid updating hidden states. During this process, the softmax function magnifies the formation of strong outliers. Building on this observation, we hypothesize that when a well-trained LLM computes gradients for text input, it often focuses these gradients on irrelevant outliers as a mechanism to prevent hidden state updates. As shown in Fig. 1, compared to SPQR, GWQ demonstrates a sparser allocation of outliers. Specifically, the outliers identified are distributed either per channel or per row, which contrasts with the more dense and uniform distribution of outliers observed in SPQR. This sparsity in GWQ results in a more reasonable allocation, as it does not require concentrating on the compensation of sensitive weight errors across every layer output. As shown in Tab. 1, with different proportions of outliers, the model using Hessian matrix search and quantized using RTN (Nagel et al., 2020) performs worse than the model using gradient search and quantized using RTN on the C4 dataset (Raffel et al., 2020), so it can be learned that it is more reasonable to use the gradient to do outliers localization.
@@ 3.1 Sensitive Weight Location
GWQ captures the absolute gradients with respect to LLM, acquired through back-propagation, the detail is shown in Alg. LABEL:gradient_alg. We disabled bias in all selected linear layers since the bias term does not get multiplied by inputs in the same way as weights do. This can cause the weight gradients to be unrepresentative of the actual importance of the weights, leading to a disproportionate influence on the MSE loss calculation. As it shown in Fig. 2, the calibration dataset is represented as $D_{c}$, the weights of LLM as $W$, and the loss function of LLM as $\mathcal{L}=(W;D_{c})$. So, the gradients $g$ of the LLM could be expressed as (2),
$g=\bigtriangledown_{W}\mathcal{L}=\bigtriangledown_{W}\mathcal{L}(W;D_{c}),$ (2)
@@
We define the quantization process as an optimization problem to search for quantitatively optimal outliers. The process can be expressed as Eq. (4),
$Q\left(W\right)=arg\min_{Q}\left\|W-W_{Q}\right\|_{2}^{2},$ (4)
Therefore the final weight optimization formula can be expressed as Eq (5),
@@ 3.2 Gradient Aware Quantization
To quantize the LLM weight $W$, we first compute the scale $s$ and the zero point $z$, which are calculated as depicted in Eq. (6) and (7), where $\beta$ represents the group size (set to 16 in this case) and $b$ is the bit-width for quantization.
@@
We select the weight by absolute values of the weight gradients $|W_{grad}|$, the largest 1% of them are filtered as outliers, and the rest are used as quantization weights. The process can be expressed as Eq. (8),
@@
The quantization and dequantization operations are carried out per channel within each group to assess the quantization error. The formulas for quantization and dequantization are provided below (Eq. (9) and (10)), and any identified outliers are excluded from the error computation Eq. (8):
$q=round(\frac{W}{s})+z,$ (9)
Finally, the quantized weight $Q(W)$ can be denoted as Eq (10):
@@
4 Experiments
Table 2: Results of Llama family on language modeling task. GWQ outperforms other quantization methods, especially when the average bit-width of the model is significantly lower.
Average Calibration Llama-2-7B-hf Llama-2-13B-hf Llama-3-8B-hf
@@ 4 Experiments
Original 16 - 5.47 / 61.01 6.97 / 56.77 4.88 / 62.88 6.47 / 57.90 6.23 / 58.34 8.99 / 51.71
+ OPTQ 4 1024 9.76 / 51.53 11.24 / 49.01 9.37/ 51.89 10.65 / 50.02 11.56 / 39.34 15.84 /44.93
+ QuIP 4 1024 8.43 / 53.13 10.01 / 50.82 8.13 / 53.48 9.49 / 51.33 10.89 / 40.84 14.03 / 46.22
+ GPTQ 4 1024 5.83 / 59.88 7.79 / 54.74 5.12 / 61.91 7.72 / 56.18 8.21 / 43.78 11.48 / 47.75
+ AWQ 4 512 5.60 / 60.54 7.70 / 55.16 4.97 / 62.53 7.10 / 56.44 6.64 / 57.07 10.69 / 47.75
+ SPQR-O 4.63 1024 5.53 / 60.80 7.02 / 56.63 4.93 / 62.71 6.50 / 57.78 6.42 / 57.70 9.23 / 51.20
+ SPQR-R 3.98 1024 5.55 / 60.56 7.06 / 55.18 4.95 / 62.55 6.53 / 56.45 6.45 / 57.09 9.26 / 48.97
+ GWQ-O (Ours) 4.63 1 5.48 / 60.87 6.99 / 56.72 4.90 / 62.76 6.48 / 53.84 6.40 / 57.83 9.11 / 51.46
+ GWQ-R (Ours) 3.98 1 5.50 / 60.59 7.01 / 56.20 4.38 / 62.56 6.50 / 56.77 6.41 / 57.69 9.15 / 49.89
Table 3: Results of QWEN-VL families on language molding task and grounding detection task. GWQ outperforms other quantization methods on both language modeling task and grounding detection task.
Average Calibration QWEN-VL QWEN-VL-chat
@@
Original 16 - 8.21 / 54.18 10.10 / 49.51 89.36 9.85 / 52.06 11.87 / 48.07 88.55
+ OPTQ 4 1024 9.34 / 53.13 12.99 /45.76 85.78 13.78 / 50.02 15.71 / 44.93 84.12
+ QuIP 4 1024 9.48 / 53.87 12.68 / 46.41 86.92 12.33 / 51.65 13.98 / 45.76 85.41
+ GPTQ 4 1024 9.31 / 52.07 11.97 / 47.08 87.13 11.50 / 52.06 13.59 / 46.09 86.67
+ AWQ 4 512 8.69 / 53.58 10.98 / 48.12 84.51 11.02 / 50.77 12.88 / 47.43 87.77
+ SPQR-O 4.63 1024 8.54 / 54.05 10.25 / 49.26 88.47 10.13 / 51.99 12.02 / 47.35 88.26
+ SPQR-R 3.98 1024 8.69 / 53.57 10.98 / 48.14 88.00 11.03 / 50.80 12.88 / 46.84 87.78
+ GWQ-O 4.63 1 8.24 / 54.06 10.10 / 49.50 88.98 9.86 / 52.03 11.88 / 48.07 88.49
+ GWQ-R 3.98 1 8.53 / 53.87 10.01 / 49.12 88.42 10.34 / 51.93 12.62 / 47.18 87.93
Overview. We first apply GWQ to the Llama2 Touvron et al. (2023b) and Llama3 Dubey et al. (2024) family of models and evaluate its feasibility on a language modeling task. We then extend our approach to multimodal models Bai et al. (2023a, b); Liu et al. (2024), measuring the generalizability of our approach to multimodal scenarios with different base models for the grounding detection task, vision language question and answer (VQA). To ensure the fairness of the experiments, we compressed the average number of bits of all the mixed-precision quantized models to less than 4 bits.
4.1 Experiments Setting
Setups. The language-only models we selected include Llama2-7B and 13B (Touvron et al., 2023b) and Llama3-8b Dubey et al. (2024). As for vision-language model, we chose LLaVA-V1.5 Liu et al. (2024), Qwen-VL Bai et al. (2023a, b), Qwen-VL-chat Bai et al. (2023a, b). These models were quantized using the first sample of the RedPajama dataset Weber et al. (2024). To evaluate Perplexity (PPL) and Accuracy (Acc.), we utilize the WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020) as validation sets. All experiments were conducted on an NVIDIA A800 (80G) GPU.